Life Insurance: A life insurance company wants to estimate the probability that a 40-year-old male will live through the next year. In a sample of 9000 such men from prior years, 8994 lived through the year. Use the relative frequency approximation to estimate the probability that a randomly selected 40-year-old male will live through the next year. Round your answer to 4 decimal places.
P(he lives through the year) =
Answer #1

No of people who lived through next year = 8994

total no of people surveyed = 9000

Hence, P(he lives through the year) = 8994/9000 = 1499/1500 = 0.9993